# Spectrum from Defocus: Fast Spectral Imaging with Chromatic Focal Stack

### v1

@@ 5 Conclusion
We have demonstrated Spectrum from Defocus (SfD), which reconstructs both RGB and hyperspectral images from chromatically aberrated focal stacks. Our camera, consisting of two off-the-shelf lenses and a grayscale sensor, achieves SOTA performance and fast reconstruction with an interpretable physics-based iterative algorithm.
Current limitations in this proof-of-concept prototype include its reliance on a specific PSF stack, limiting its performance to our predicted depth range and a paraxial field of view. Expanding beyond our current working volume will require a combination of both optical and algorithmic engineering, to ensure that signals of sufficient strength are available and can be recovered accurately and efficiently. Additionally, because we optimize our method for reconstruction of a high dimensional spectrum, we require $>3$ measurements for RGB images. Though we ensured a fair comparison on static scenes by giving all methods the same total exposure time in simulation, an important future direction for dynamic imaging is to bring down the number of measurements needed through a combination of lower-dimensional output targets with a heavier use of training data.

### v2

@@ 5 Conclusion
We have demonstrated Spectrum from Defocus (SfD), which reconstructs hyperspectral images from chromatically aberrated focal stacks. Our camera, consisting of two off-the-shelf lenses and a grayscale sensor, achieves SOTA performance and fast reconstruction with an interpretable physics- and data-based iterative algorithm. We demonstrate its performance in simulation against eight other methods, with functional principles including spectral filters, dispersive elements, diffractive optical elements, and chromatic focal stacking, apried with reconstruction algorithms based on physical models, data-driven networks, and hybrid approaches. Our hardware prototype matches simulation performance on real-world data.
Our proof-of-concept prototype also reveals several limitations. First, we assume scenes occupy a working range of 34 cm (see supplement, but note robustness beyond theoretical predictions demonstrated in Fig. 5) and a narrow paraxial field of view (see supplement, but note comparable limitations on almost all SOTA methods in Tab. 1). Expanding the working volume will require joint optical and algorithmic design to ensure that signals remain strong and recoverable across a broader range of conditions. Second, red and black colors experience slightly lower reconstruction accuracy. This is because our chromatic focal shift flattens at longer wavelengths, which could be improved with a more dispersive optical glass, and because the system exhibits limited dynamic range. Rare failure cases tend to occur in black regions, which can appear gray or falsely colored due to sensor saturation and light bleed from neighboring bright pixels. Third, because the reconstruction must model incoming photons of all wavelengths, a spectral reconstruction is required in order to produce RGB measurements. In practice, this means that five measurements are required to produce high-quality RGB images. Reducing the effects of multi-shot imaging, using tools from burst photography [14, 20, 30] and greater reliance on data-driven reconstruction, is an important direction for SfD in dynamic scenes.
Finally, while the light efficiency and robust reconstruction of our method are well-suited to fast hyperspectral imaging, and preliminary results on low exposure times suggests a significant advantage to our method in photon-starved settings (see Fig. 3b and supplement), exploring this advantage will require higher fidelity noise models for low-light images [40, 36].